EVA Report – April 3rd

Crew 279 EVA Report 03-04-2023
EVA #01
Author: Ioana Dimitrova
Purpose of EVA: Training EVA + communication test with crew members and with the hab.
Start time: 9:20
End time: 10:10
Narrative: The goal of the EVA was to get comfortable with the spacesuits, the protocole and the environment. Preparation went well but during the decompression in the airlock, we had trouble hearing what one of the crewmembers was saying. We aborted the decompression and tried to fix the problem. We put the walkie talkie on the other side of the person’s chest, played with the cable a little bit, redid the whole procedure for radio and helmet and tried again, everything seemed to be working fine. The first part of the EVA consisted of getting to the Marble Ritual with the rovers. The second part was to walk around the Marble ritual area to discover the place and to get acquainted with the spacesuits, the radio communication and the particularities of the soil. This walk consisted of 20 min of the EVA. No communication problems seemed to exist. Exploration was nominal. The third part was the return to the Hab with the rovers. Upon arrival, the crewmember who had a problem with the radio prior to the mission, tried to communicate the rover numbers to the Hab and it was at that moment that we noticed that his comms were not working properly. It was time to end the mission anyways, so we went inside after decompression.
The crewmember in question is not the chatty type, so even if the problem occurred earlier in the EVA, we didn’t notice the problem then. Our hypothesis is that there is a problem with the headset because we tested the radio with another headset and it seemed fine. The headset doesn’t seem to be able to be fixed. During the next EVA, the crewmember in question will use another headset with his radio.
Destination: Marble Ritual
Coordinates (use UTM WGS 84): 519000 E ,4250500 N
Participants: 4: Ioana (Engineer) (EVA Leader), Agnes (Executive officer), Gwen (Astronomer), Augustin (Greenhab officer).
Road(s) and routes per MDRS Map: Cow Dung Road 0110
Mode of travel: First drive of Rover with spacesuits and accommodation with maneuver to the Marble Ritual then walk around no farther than 200 meters for communication check.

Operations Report – April 3rd

Crew 279 Operations Report 03-04-2023
SOL: 1
Name of person filing report: Ioana Dimitrova
Non-nominal systems: robotic observatory
Notes on non-nominal systems: awaiting for a solution from the manufacturer
Spirit rover used: no
Beginning charge: NOT MEASURED
Ending charge: NOT MEASURED
Currently charging: YES
Opportunity rover used: no
Beginning charge: NOT MEASURED
Ending charge: NOT MEASURED
Currently charging: YES
Curiosity rover used: yes
Hours: 226.8
Beginning charge: 100%
Ending charge: 92%
Currently charging: yes
Perseverance rover used: yes
Hours: 260.1
Beginning charge: 100%
Ending charge: 91%
Currently charging: yes
General notes on rovers: Spirit and Opportunity not used today and crew forgot to measure them but will be measured tomorrow and the days to come.
Summary of Hab operations: Music influence on stress experiment was conducted. Inventory of food and what everyone ate today for microbiota experiment. Knobs on oven fix (see below for more info).
WATER USE: (see below in questions for mission support) Will be measured tomorrow
Water (static tank): 484.65 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 14:00 to 21:00 for experiments and reports
Summary of suits and radios:
Suits nominal, charging nominal. Radios nominal but technical problem with headset number 5. I think the problem is in the cable of the microphone. Sometimes it works, sometimes not. From other radios, we can always hear that radio number 5 is transmitting something (with the click sound of a walkie talkie), but often you cannot hear the person at all. Which leads us to think the problem is not with the radio itself. We also tried changing headsets with radio number 5, and there is no problem. So the problem must be in the headset, and because sometimes it works and sometimes not, I suppose it’s a cable problem.
We noticed this problem at the beginning of the EVA in the airlock. We aborted the decompression and resolved the problem before decompressing again and going outside. Augustin (who was wearing the headset in question) is a very quiet person, he did not talk much during the EVA, that is why we noticed the problem only when he tried to communicate the rover numbers to Hab at the end of the EVA when we were about to enter the Hab.
Solution: The problem does not look like it could be resolved apart from replacing the cable or the headset entirely. When radio number 5 has to go out on EVA, we will give him another headset to use.
Summary of GreenHab operations:
WATER USE: 13.4 gallons
Heater: Off
Supplemental light: Off
Harvest: cherry tomatoes 217 g, Basilic 4g, Mint 3g, Sage 0.5g
Summary of ScienceDome operations: extremophiles were put in cultures and incubated overnight. Arduino was programmed for pressure, altitude, temperature and humidity for a weather balloon.
Dual split: Heat or AC, Off
Summary of RAM operations: Assembly of 3D printer. Knobs on the oven were complicated to use so we worked a little bit on them and they work better now.
Summary of any observatory issues: see above
Summary of health and safety issues: Hole in the tunnel in front of the Science Dome because of wind gusts. Commander was blocked in there during repairs. Repairs were made by Sergii, everything nominal now.
Questions, concerns and requests to Mission Support:
Concerning water use, the water tank was filled today to 520 gallons during the day, and I measured the water level at the end of the day but not before the fill.

GreenHab Report – April 3rd

Crew 279 GreenHab Report 03-04-2023

GreenHab Officer: Augustin Tribolet
Environmental control: Open during day and heater during night
Average temperatures: 79.7 F
Hours of supplemental light: N/A
Daily water usage for crops: 13.4 gallons
Daily water usage for research and/or other purposes: N/A
Water in Blue Tank: 172.8 gallons
Time(s) of watering for crops: 10:30 am, 17:20
Changes to crops: N/A
Narrative: As part of the day-to-day activities of the GreenHab, general watering and maintenance of plants and crops was done, no new planting was performed. The door was left closed most of the day due to strong wind. A mouse has been caught today and released in the wild far from the station. The mouse trap will be reinstalled this evening.
Harvest: 217 g tomatoes, 4 g basilic, 0.5 g sage, 3 g mantes
Support/supplies needed: Watering can made of tin has a hole in it, but plastic watering can is sufficient

Journalist Report – April 3rd

Crew 279 Journalist Report 03APR2023

Journalist report : Aglaé Sacré

SOL1 : Arrival on Mars

Last night, our ship finally arrived on the planet Mars.
After a first visit to our Crew Health and Safety to fill a questionnaire about our feelings this morning and an analysis of our sleep via an application, we had our first Martian breakfast before leaving for the first spacewalk.

Despite the wind, half of the team went for a first outing to test the rover and to acclimatize to the suits and the spacesuit but also to learn how to handle the talkie which allows the communication between the station and the crew members outside but also between the crew members during the outing. They went to the Marble ritual.

After an hour of walking in the surroundings of the station, the 4 astronauts returned, and it was only after a five-minute depressurization that they could enter the station.

During this time, the other part of the team has been working on the inventory of the food present in the station to establish the menus and the quantities for the next two weeks.

It’s already time for our first dinner in the station, Ttele and Aglaé are the first to go behind the stove. On the menu of the day, rice with freeze-dried vegetables, a real success for the first time.

During the afternoon, we could start working on our personal experiences.
Our Green Hab Officer harvested the first tomatoes in the greenhouse that we will eat tonight. We also built a 3D printer that will be used for some experiments.

The experiment on the influence of music on stress has also started, a part of the team spent a quiet moment with their favorite music to calculate the impact of this music on their stress.

At 7pm it is already time to send all our reports to the base to ensure the smooth running and the good maintenance of the station. We are also planning tomorrow’s spacewalks.

On the menu tonight, sweet potato and vegetables!

Conclusion of this day: a first day rich in discoveries, our experiments are starting, and we are eager to see how they will evolve.

Sol Summary – April 3rd

Crew 279 Sol Summary Report 4-3-23

Sol: 1

Summary Title: A windy and intense start

Author’s name: Antoine de Barquin – Crew Commander

Mission Status: Nominal
Sol Activity Summary:
First awakening for the Ares crew on the Martian soil. The night was short and turbulent because of the wind but the team started the day more motivated than ever. We are now occupying this planet and intend to discover all of its secrets. The first step on the ground came at 9:15 am during the first EVA training. The team was able to practice moving in their suits, driving the rovers and communicating with each other and with the base.
During this time, the rest of the crew could realize an inventory of the material and the food in order to organize the continuation of the simulation.
The work of the previous teams was appreciated during the tasting of the first tomatoes harvested in the GreenHab ! The end of lunch launched the start of each crew member’s experiments. While some built a 3D printer in the Ram, others developed a bacterial culture in the Science Dome.
At the end of the day the strong winds caused a breach in the tunnel blocking the commander in the science dome for the time of the repairs.
Look Ahead Plan:
In Sol 2 a second EVA training is planned for the rest of the crew. This EVA will finish the training of the crew and the outdoor experiences will start on Sol 3.

Anomalies in work:
Headset malfunction to be repaired tomorrow
Breach in the tunnel leading to the science dome. Repair already done

Weather: Very windy, high 16°C/61°F, low 5°C/41F°

Crew Physical Status: Optimal

EVA: See above

Reports to be file: Sol Summary, Operations Report, Journalist Report, HSO Report, GreenHab Report

Support Requested: None at this time.

Copyright © The Mars Society. All rights reserved. | Main Site